OpenStudy (anonymous):

can someone help ? Hailey and Matt paddled a canoe from a dock to their campsite. It took them 3 hours to get from the dock to the campsite while traveling with a 3 mph current in the water. It took them 5 hours to return to the dock against the same current. What was their speed in still water? a. 10 mph b.12 mph c.14 mph d.16 mph

OpenStudy (studygurl14):

This should help: http://voices.yahoo.com/algebra-help-solve-boat-river-word-problems-2820345.html :)

OpenStudy (anonymous):

r t = d Let r be the speed in still water. 3 (r + 3) =5 (r - 3) r = 12 mph

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (studygurl14):

you're welcome

OpenStudy (anonymous):

if x was their speed, they were going x+3 for 3 hours there. they were going x-3 for 5 hours. 3(x+3)=5(x-3) 3x+9=5x-15 9+15=5x-3x 24=2x x=12